### v1

# Aptly: Making Mobile Apps from Natural Language

Abstract
We present Aptly, an extension of the MIT App Inventor platform enabling mobile app development via natural language powered by code-generating large language models (LLMs). Aptly complements App Inventor’s block language with a text language designed to allow visual code generation via text-based LLMs. We detail the technical aspects of how the Aptly server integrates LLMs with a realtime collaboration function to facilitate the automated creation and editing of mobile apps given user instructions. The paper concludes with insights from a study of a pilot implementation involving high school students, which examines Aptly’s practicality and user experience. The findings underscore Aptly’s potential as a tool that democratizes app development and fosters technological creativity.
Index Terms:
MIT App Inventor, Large Language Model, Block programming
§§footnotetext: Equal contribution
I Introduction
The rise of technology in daily life highlights the need for everyone, particularly youth, to engage with digital tools to solve real-world challenges. Mobile applications serve as a powerful medium for such engagement given the widespread adoption of smartphones and tablets across diverse demographics, including under-represented communities [1]. However, the traditional app development process, which often requires a solid foundation in computer science, has historically excluded many aspiring creators from participating. To democratize this process, the MIT App Inventor initiative has revolutionized the creation of mobile apps by introducing a user-friendly, visual programming environment that empowers anyone, regardless of their technical background, to develop functional applications [2]. App Inventor utilizes a block-based programming approach, which allows users to visually assemble their applications using geometric shapes, making the development process more intuitive and accessible [3]. The tool not only fosters creativity among novice programmers but also encourages a more inclusive approach to technological education and development [4]. Tissenbaum et al. [5] call the process by which people leverage technology like App Inventor to make societal impact computational action.
The advent of large language models (LLMs) and their ability to generate code [6, 7] has opened new avenues for mobile app development. Capitalizing on this technological advancement, our research introduces Aptly, a groundbreaking platform that allows users to create mobile applications using natural language to lower barriers to computational action. For example, a user could simply state, “Make me an app that translates English to Spanish,” and Aptly would automatically generate a functional app based on this input. This paper details the framework of Aptly, beginning with an overarching view of the system’s architecture. We will discuss the development of the Aptly Language and our method of integrating it with an LLM to produce the desired syntax. Additionally, we explore the mechanisms that enable users to refine their apps through natural language commands. We conclude by examining Aptly’s current functionalities and presenting preliminary findings from its deployment among high school students.
II Aptly Architecture
Figure 1: An interaction diagram between the different components of the MIT App Inventor with Aptly system.
Aptly augments MIT App Inventor with generative AI coding capabilities. We accomplish this by building on a version of App Inventor supporting real-time collaboration (RTC) [8]. The Aptly agent acts as a collaborator with users, taking their instructions in natural language and manipulating the project just as another user might. The overall system involves the following components:
  • MIT App Inventor: The base version of the App Inventor [^1] https://github.com/mit-cml/appinventor-sources is extended to integrate with the Aptly system and the RTC server to facilitate enhanced functionality.
  • MIT App Inventor RTC: The real-time collaboration (RTC) server mediates between agents in the system architecture.
  • Aptly Server: Detailed in the subsequent section, the Aptly server invokes the LLM and converts its responses into events understood by App Inventor RTC.
  • OpenAI GPT-3.5/4: Utilizing OpenAI’s LLM models, this component provides advanced code completion services [9, 10].
Figure 1 depicts the relationship between the various components as they are used by the app creator.
II-A The Aptly Language
Screen1 = Screen()
HA1 = HorizontalArrangement(Screen1)
Label1 = Label(HA1, Text = "Weight in lbs: ")
EarthWeight = TextBox(HA1, NumbersOnly = True)
Label2 = Label(Screen1, Text = "Select Planet:")
PlanetList = ListView(Screen1, ElementsFromString = "Mercury, Venus, Mars, Jupiter, Saturn, Uranus, Neptune")
Calculate = Button(Screen1, Text = Calculate’)
PlanetaryWeight = Label(Screen1)
initialize gravities = {"Mercury": 0.38, "Venus": 0.91, "Mars": 0.38, "Jupiter": 2.34, "Saturn": 0.93, "Uranus": 0.92, "Neptune": 1.12}
to compute_weight(earth_lbs, planet):
return earth_lbs * dictionaries_lookup(planet, global gravities, not found’)
when Calculate.Click():
set PlanetaryWeight.Text = call compute_weight(EarthWeight.Text, PlanetList.Selection)
Listing 1: An app to calculate weights on different planets expressed in Aptly.
Figure 2: An example of the Aptly code in Listing 1 and its correspondence to the App Inventor interface.
The Aptly language is designed to be used with LLMs. It is a textual representation of App Inventor programs inspired by Python. The decision to base Aptly on Python was made for two reasons:
  1. 1.
    Python’s pseudocode-like syntax closely aligns with the readability and structure of App Inventor blocks, facilitating intuitive understanding and use.
  2. 2.
    LLMs have been extensively trained on Python code, enhancing their ability to generate accurate and functional code in a similar syntax.
Given these attributes, we hypothesize that generative models can be effectively fine-tuned to interface with Aptly, leveraging its Python-like structure. However, it is important to note that while Aptly draws inspiration from Python, it is distinct. Aptly has been specifically constrained to ensure a one-to-one correspondence with App Inventor: every valid App Inventor program is also a valid Aptly program, and vice versa. See Listing 1 for an example Aptly program and Figure 2 for showcasing how the code is represented as blocks.
Figure 3: When the user requests an application with its textual description, we automatically synthesize the prompt by adding several example pairs along with the desired application’s textual description. This constructed prompt is fed into one of OpenAI’s GPT-X models as an input, which outputs code that can be converted into a fully functional mobile application
To ensure the fidelity and accuracy of the one-to-one correspondence between the Aptly and App Inventor languages, we employ a comprehensive suite of integration tests. These tests validate the conversion cycle by processing Aptly code through a round-trip transformation to an App Inventor (.aia) file and back to Aptly. Furthermore, any new additions to the Aptly language introduced through pull requests are required to include corresponding tests that affirm the functionality and behavior of the new features.
For maintaining the precise correspondence of the Aptly language, we define the following four components:
  • Parser: Parses the Aptly code into its respective representation in the Abstract Syntax Tree (AST). Each method adheres to the extended Backus-Naur Form (eBNF) of the Aptly language.
  • ProjectReader: Transforms the App Inventor representations into the corresponding Aptly AST.
  • ProjectWriter: Converts Aptly AST into the App Inventor representation (either in the designer or blocks format).
  • Serializer: Outputs the AST into the Aptly code format.
In tandem, these four components provide the core functionality to enable app creation with Aptly by allowing us to convert between the text and block representations via the Aptly AST.
Figure 4: An app created from a prompt given in Dutch.
II-B Few-shot Prompt Engineering
Aptly leverages machine learning research on generative models, which are probabilistic frameworks that learn how datasets are generated to facilitate new data creation. We specifically employ a transformer-based large language model to develop mobile apps [11]. This approach capitalizes on the recent significant advancements in large language models, an area that has garnered substantial interest for its impressive research outcomes and substantial commercial opportunities. Generative models operate based on “prompts” which are crucial inputs that significantly influence their performance. The effectiveness of these models hinges on the quality of the prompts, a practice known as prompt engineering—an active field within machine learning research [12]. A common method of prompt engineering involves “few-shot” prompts, where a model is given a limited number of solved task examples as part of its input. This approach is often interpreted as the model “learning” the task in real-time from these few examples [9]. Aptly uses Open AI’s LLMs, which are capable of generating computer code from natural language descriptions in several programming languages including Java, JavaScript, Python, C, C#, and Swift.
When the user requests an app with their natural description, we synthesize a prompt which is a natural language description (denoted as $D$) of the desired app to be created, together with a set of example pairs, such as the following $<<d\_1,c\_1>><<d\_2,c\_2>>...<<d\_k,c\_k>>$, where $d\_i$ is the description of application $i$, and $c\_i$ is the corresponding Aptly code of the description for application $i$. The example pairs come from a database of unique Aptly examples compiled by the team from apps created on the App Inventor platform. The example pairs are not expected to include the literal description $D$ to be processed nor the actual Aptly Code to be generated which would make the generation task trivial. Instead, generative models use the provided examples to guide their processing in generating new original output.
Aptly’s method for providing relevant prompts to accompany a description $D$ is to provide a set of example application descriptions paired with appropriate Aptly Code. We use semantic similarity to highlight the “appropriateness” of code and description. The method relies on having a way to automatically measure the similarity of text descriptions and the program. For computing the similarity of the text description and program, we use embeddings of data elements, which are numerical representations of concepts converted to number sequences [13]. In our scenario, an embedding represents the semantic meaning of a natural language description or code. To measure the similarity of two items, one embeds them in the same vector space and takes the distance between the two embedding vectors as a measure of similarity (Small distance implies high similarity while large distance implies low similarity). We use cosine distance, which reflects the angle between vectors, to compute the similarity between two vectors.
We employ OpenAI’s Text/Code Embedding model to generate embeddings from descriptions and associated code. These neural network models utilize Contrastive Pre-Training—a training method that functions by clustering in vector space [14]. Specifically, it draws predefined positive examples (i.e., matching text and code) closer together, while pushing negative examples (i.e., contrasting text and code) apart. This technique enhances our model’s ability to identify and select the example pair most relevant to a user’s specified description. We feed the synthesized prompt into GPT, which outputs the Aptly Code corresponding to the user’s requested app description. We can then convert the generated code into App Inventor blocks to generate a fully functional application (Figure 3).
An interesting feature of the use of LLMs for this process is that the user input need not be expressed in English as might be the case with traditional natural language processing techniques. As an example, we provided this prompt in Dutch: “Maak een app met een tekstbox, een lijst met zes talen en een knop met de tekst ‘vertalen’. Wanneer er op de knop wordt gedrukt dan vertaal je de tekst in de geselecteerde taal en laat de vertaling zien.” This produced an app that gave the user the option to translate into English, Spanish, French, German, Italian, and Japanese, with all of the app content expressed in Dutch (Figure 4). Importantly, all of the example pairs we provide the LLM are written in English–there are no examples expressed in Dutch. Ultimately, these abilities of LLMs will reduce the barrier to computational action by removing the need to learn English before or while learning to code.
Figure 5: An overview of the information flow during an edit operation using Aptly. Arrows represent the flow of information and should be read as “tail label head,” e.g., “App Inventor sends AIA file” or “RTC Event Sequence consumed by App Inventor RTC.”
II-C The Editing Capability
Often it is unreasonable to assume one can create a full app in a single shot request. We aim to empower programmers not only to generate apps using natural language, but also to refine, iterate, and evolve their designs continuously through natural language commands. This approach allows for a more dynamic development process, enabling creators to adapt and modify their applications easily as new ideas or requirements emerge. This section details how editing of apps is achieved in Aptly.
Figure 6: The update operations (blue) to transform the AST of the original program into the AST of the modified program. Some unmodified nodes are omitted for brevity.
II-C1 Editing Projects using Descriptions
Editing a project with Aptly is done by emitting a sequence of edit operations over the MIT App Inventor RTC system. When a user requests an edit by providing a description of their desired change, App Inventor sends the AIA for the current project plus the user-provided edit description to the endpoint of the Aptly server.
The editing process uses the following steps (see also Figure 5). The Aptly frontend sends the project state and user request to the server (1), which utilizes the ProjectReader and Serializer to obtain the corresponding Aptly code. The server then constructs an edit prompt–a top-ranked selection of example edits plus the project (converted to Aptly code) and the description–to be sent (2) to the LLM. We use the same technique for ranking examples as discussed previously.
When the LLM returns updated Aptly code, the Aptly server constructs the AST representation of the original code (via the ProjectReader) and the modified code (via the Parser), and computes the minimum-cost sequence of edit operations to transform the original AST into the modified one. The Aptly server utilizes the tree edit distance algorithm proposed by Zhang and Shasha [15] to compute the minimum-cost sequence of edit operations (Fig. 5 note 3). Aptly specifically utilizes the Python zss module.The edit operations returned are INSERT, REMOVE, UPDATE, and MATCH. The update_cost function is customized for App Inventor by taking into account the type of two nodes; for example, if there are two component declarations (e.g., Button1 = Button(...) and Label1 = Label(...)), there should be infinite update cost to transform one of those component declarations into the other because Button and Label are two different types and App Inventor does not allow a way to change a button into a label or vise versa. Similarly, in App Inventor, if there are two bound component events (e.g., when Button1.Click() and when Button2.Click()), it should only be possible to update one to be the other if the type of the component is the same because this is the only type of change App Inventor allows in the Blocks editor for a bound component event. If the user wants to change which method they’re using (e.g., Click vs. TouchUp) or change the event to a different component type (e.g., Button1.GotFocus vs Switch1.GotFocus), the user would have to grab a new block from the blocks drawer and delete the old block, so Aptly cannot make this change via an UPDATE operation either.
Given the sequence of insertions, removals, updates, and matches recommended by zss, the TreeDiff class emits a sequence of edit events compatible with the App Inventor RTC system (Fig. 5 note 4). These edit events include creating, updating, moving, and deleting components and blocks. Some changes require multiple edit events; for example, since block stacks are represented as linked lists, inserting a block in the middle of a block stack requires creating the block, moving it to the correct location (i.e., setting its parent correctly), and moving its descendent (if applicable) to now point to it rather than its previous parent. As another example, deleting an item in a list or deleting a pair in a dictionary requires issuing a move event for all items after the deleted item in the list, as even though these structures are not represented as linked lists, each subsequent item’s input name (e.g., ‘ITEM1’ vs. ‘ITEM0’ for list items) will change.
To emit the correct edit events via the RTC system, we process the operations returned from zss in several stages:
  • Preinsertion: To handle moving components, we need to perform a preprocessing step that aggregates all of the new components into a map based on their name and type. We process all of the insert operations to build this map, which will be consumed as part of the remove stage to determine whether a component was renamed, moved, or removed. If we did not include such a stage, we might end up deleting and then inserting the same node rather than moving a node, incurring unnecessary cost.
  • Remove: We emit removal events for deleted components first. When a component is removed, we log this information so that any block removals involving that component are ignored since they are covered by the component’s removal. Likewise, we need to log all of the block IDs that are removed so that we do not also emit removal events for the subtree starting at that node. Note that there is a special case in regards to moved components; for subtrees starting at nodes that are marked as insertions, a removal does not take place.
  • Match: The Match operation in the tree edit sequence is a no-op. However, when translating the operations into App Inventor events, we use this step to perform some bookkeeping necessary to support insertion operations later.
  • Update: The Update stage converts the value of a node in the source tree to a new value at the same position in the destination tree. Common reasons for update operations involve renaming components, changing property values, or renaming variables or procedures.
  • Insert Stage: The Insert operation adds a new subtree to the AST. Because we enumerate the changes in breadth-first order, an insertion in the tree covers the whole subtree at that node. When subtrees are inserted, we must return to looking at the parent node (which would be a match/update operation) in order to issue the necessary move events needed based on the inserted node (such as moving the new node to the right place and adjusting the parent and input name of successors if needed).
Once all of the edit events have been created, the Aptly server joins the user’s project in the RTC system and emits the stream of changes (Fig. 5 notes 5, 6).
II-C2 Applying Updates via Real-Time Collaboration System
Aptly leverages the App Inventor Real-Time Collaboration (RTC) [8] system to edit the project based on the event sequence produced in the previous step. Aptly is conceptualized as a collaborative user within the project environment. Once the event sequence is generated by the Aptly server, it joins the user’s project in the RTC server. It then sends the event sequence to the server, which relays the changes to the user’s App Inventor session to be applied to the project. Leveraging the RTC system, Aptly allows for modifications to be reversed if they do not meet the user’s expectations. This functionality is facilitated by the RTC’s capability to execute events in reverse, relying solely on a stack of events that can be undone, thereby eliminating the need for maintaining additional state information. This design enhances user control and flexibility, making the editing process more user-friendly and adaptable to changes.
II-C3 Worked Example
Consider the following Aptly code representing the app “HelloPurr,” App Inventor’s equivalent to a Hello World program.
Screen1 = Screen(AppName = "HelloPurr")
Cat = Button(Screen1, Image = "kitty.png")
Meow = Sound(Screen1, Source = "meow.mp3")
when Cat.Click():
call Meow.Play()
One might offer in natural language “add a label below the kitty that says ‘pet the kitty’.” The LLM may generate the following Aptly code in response:
Screen1 = Screen(AppName = "HelloPurr")
Cat = Button(Screen1, Image = "kitty.png")
Label1 = Label(Screen1, Text = "pet the kitty")
Meow = Sound(Screen1, Source = "meow.mp3")
when Cat.Click():
call Meow.Play()
An illustration of the above two codes AST, respectively noted as $a$ and $b$, is shown in figure 6. The two ASTs are then processed into a sequence of edits needed to transform the AST of the first snippet into the AST of the second snippet. In this particular scenario, the Zhang-Shasha algorithm will result in MATCH operations except for the INSERT operations that introduce the AST nodes corresponding to the insertion on line 3:
INSERT Label1 (Identifier)
INSERT Label (Reference)
INSERT Screen1 (Reference)
INSERT Text (Reference)
INSERT "pet the kitty" (StringValue)
INSERT Text = "pet the kitty" (DesignerAssignment)
INSERT ComponentTypeInvocation(Label1, Screen1, {Text: "pet the kitty"}) (ComponentTypeInvocation)
INSERT ComponentDecl(Label1, ComponentTypeInvocation(Label1, Screen1, {Text: "pet the kitty"})) (ComponentDecl)
For the processing of INSERT operations, Aptly considers the operations in a breadth-first fashion, which leads to the ComponentDecl being the first node, n, encountered during traversal. We know that an insertion of a node a depth d subsumes all of the children below depth d in the subtree rooted at n. The TreeDiff class checks to see if it has a method to handle the insertion of the AST type of n (i.e., ComponentDecl). If such a method exists, it is called. Otherwise, the subtree starting at n will continue to be traversed breadth-first. In the case of ComponentDecl, such a method does exist and it interprets the insertion of a ComponentDecl node as the RTC event sequence of ComponentAdd, ComponentMove, and one or more ComponentProperty events. For example, if the UUID of the label component is 12345:
ComponentAdd(12345, "Label", [])
ComponentMove(12345, 0, 1)
ComponentProperty(12345, "\$Name", "Label1")
ComponentProperty(12345, "Text", "pet the kitty")
This sequence of events is then streamed to the user’s workspace via the RTC server.
III Results
III-A Coverage of App Inventor
Table I summarizes the current coverage Aptly provides for App Inventor functionality. We achieved our target of having one to one correspondence of the block language, with a few exceptions for new language features added by the App Inventor team since the Aptly project began.
TABLE I: Summary of Aptly’s block language coverage
Category Coverage %
Controls 100
Logic 100
Math 100
Text 94
Lists 100
Dictionaries 94
Colors 100
Procedures 75
Variables 80
Components 80
Total 96
TABLE II: Summary of Aptly Examples’ coverage of App Inventor functionality.
Feature Coverage %
Components 40.0
     Events 13.6
     Methods 6.4
     Properties 13.7
Coverage of App Inventor’s component library is more limited as shown in Table II. When building examples, we have aimed to cover the most common functionality observed in projects as studied by the App Inventor team [16]. Importantly, there is a very long tail of component usage that suggests that our initial set of example projects likely covers most use cases encountered by App Inventor users.
III-B Pilot User Study
To evaluate the effectiveness and user experience of our tool, Aptly, we conducted a study involving 10 high school students with the following demographics:
  • Age: 17 (5), 18 (5)
  • Gender: Female (8), Male (2)
  • General programming experience: None (3), $<$ 1 year (2), 3-4 years (3), 5+ years (2)
  • Familiarity with block-based programming: None (3), Beginner (2), Familiar (5)
This diverse group was selected to yield insights into how students with different expertise levels and app creation familiarity engage with our tool. The study duration ranged from 50 to 80 minutes per participant. Initially, we introduced the students to Aptly, demonstrating how to create a basic app with interactive features such as a button and a label, and how to modify properties like color and font using the tool’s editing capabilities. We then assigned two distinct tasks to assess the tool’s utility and the participants’ ability to leverage it.
Figure 7: App Inventor Basic Tutorial: Digital Doodle
III-B1 Recreation Task
Students were shown “Digital Doodle” (Figure 7), a beginner-level app from App Inventor, and asked to recreate it using Aptly. This task was designed to observe whether participants could effectively translate a clear visual concept into a functional app through verbal descriptions, reflecting their understanding and communication of app design concepts. The task demonstrated the platform’s flexibility in understanding and executing a range of instructions, from detailed descriptions specifying colors and functionalities of buttons to succinct commands like “Draws with RGB.” This versatility highlights Aptly’s capacity to interpret and implement user input effectively, allowing users with no expert programming skills to successfully create functional apps.
Participants employed different strategies in app creation, some opting for a more incremental approach using Aptly’s editing features. For instance, one participant began with a basic layout and progressively added functionalities and aesthetic adjustments, such as alignment and color changes, through subsequent commands. This iterative process was particularly favored by participants with programming backgrounds, reflecting a workflow similar to traditional software development where features are built and refined over time. This suggests that Aptly can accommodate various user preferences and workflows, enhancing its usability across different skill levels and development styles. However, the study also uncovered challenges related to the precision of language used in commands. Some participants struggled to articulate their desired app functionalities clearly, leading to discrepancies between their intentions and the app’s behavior. For example, a participant’s command to “draw a line when I click a button” was too vague without specifying the color interaction, illustrating the current limitations of AI in interpreting user intent without explicit instruction.
III-B2 Scenario-Based Task
Participants were given a scenario. Specifically, they were asked the following scenario: “In your class you have a friend that is visually impaired and is having trouble with their math class. You want to help that student by creating a calculator app that will help them with their addition and subtraction. Try using Aptly to create that app. Feel free to use the editing and also manually drag and drop the blocks.” This task aimed to evaluate participants’ ability to ideate and articulate solutions to real-world problems using Aptly. It also encouraged them to consider accessibility features, such as voice commands and easy-to-navigate interfaces, enhancing their problem-solving and design-thinking skills.
Participants provided initial descriptions that ranged from simple one-liners like “Make an app that is a calculator” to detailed paragraphs specifying the layout and functionality, including voice recognition and auditory feedback. This task demonstrated the ability of participants to conceptualize and articulate diverse solutions using Aptly, catering to a specific accessibility need. Participants’ approaches varied; some chose to incrementally build and refine their app’s user interface and functionalities using Aptly’s editing tools, while others outlined their complete vision from the onset. For example, one participant started with a basic calculator design and progressively enhanced its features and interface to be more accessible for visually impaired users. This approach highlights Aptly’s flexibility to accommodate different development styles and the iterative nature of designing user-centric applications.
Moreover, the participants’ creative solutions for accessibility features showcased the potential of integrating advanced technologies such as text-to-speech and tactile feedback into apps. One innovative approach involved using screen taps and phone vibrations to communicate numbers and results, demonstrating how users can leverage Aptly not only as a development tool but also as a platform for ideation. The diversity in usage and the degree of reliance on Aptly’s capabilities varied, indicating that while some users fully embraced the AI-driven development process, others preferred a more hands-on approach in shaping their applications. This underscores the importance of versatile tools like Aptly in modern app development, especially when addressing the specific needs of diverse users.
III-B3 General Observation
Observations revealed valuable insights into user behavior and platform design. Participants exhibited a general impatience with the system’s response times, often proceeding to plan subsequent changes if immediate implementation of edits was not observed. Challenges arose when Aptly misunderstood user commands; while some participants attempted to rephrase their inputs multiple times using Aptly editing, others opted to manually make the changes. Notably, more experienced programmers frequently preferred manual adjustments and issued a higher number of commands compared to their less experienced counterparts. These seasoned programmers expressed a desire for greater control, including access to error messages to better understand the system’s handling of their commands. This feedback highlights a tension between designing intuitive interfaces suitable for beginners and providing detailed control and feedback that experienced users seek. This dichotomy suggests that while simplifying interfaces can benefit novices, it may frustrate advanced users who expect deeper interaction and transparency from the development tools.
III-B4 Participant responses
To capture the changes in participants’ perceptions before and after using Aptly, we asked the participants two questions before and after the experiment.
  • “Programming is accessible for everyone”: Prior to the workshop, students’ views on programming accessibility varied, with some expressing skepticism about universal accessibility due to factors such as technological availability and socioeconomic barriers. Common responses ranged from “somewhat disagree” to acknowledging partial agreement, contingent on access to technology like computers and the internet. After using Aptly, participants generally showed a nuanced view, with some noting improvements in their perceptions of accessibility. Notably, though some recognized improvements, they still highlighted limitations, particularly for individuals who might lack basic digital literacy or access to necessary hardware.
  • “Artificial Intelligence can be a powerful tool to help people in programming”: Initially, responses indicated a positive sentiment towards AI as a powerful aid in programming, with a slight variety hinting at potential job displacement concerns. After the workshop, reflections were more detailed, emphasizing the practical benefits and challenges observed during the use of AI-driven tools like Aptly. Participants appreciated AI’s potential to simplify the programming process, especially for beginners, yet also discussed the importance of understanding underlying logic beyond AI assistance. This shift suggests a more informed appreciation of AI’s capabilities and limitations in programming contexts.
We also asked each participant the question “Did Aptly change your interest in learning how to program?” on a scale of 1 to 5, where higher values signify increased interest. The range of responses spanned from a low of 2 to a high of 4, with no extreme sentiments expressed, as the ratings did not include the lowest (1) or highest (5) possible marks. This distribution implies that while Aptly was generally effective in increasing interest among most participants, it had a varied impact. The mean response was 3.7 (SD = 0.67). The mixed response highlights the potential of Aptly to engage users, albeit not uniformly, suggesting avenues for further investigation into individual differences and tool optimization to broaden its appeal and effectiveness.
III-B5 Limitation of Aptly based on responses
In our study, participants identified several limitations of the Aptly platform, providing valuable insights into its operational shortcomings and areas ripe for development. Notably, Aptly struggled with complex command executions such as playing specific audio files, drawing predefined shapes, and making API calls. Furthermore, it demonstrated limitations in processing abstract or intricate commands, which underscores deficiencies in its natural language processing capabilities.
Feedback from users on improving Aptly emphasized enhancing both functionality and user interface design. Participants suggested increasing the accuracy and range of command interpretation to reduce errors, thereby making the platform more robust for complex programming tasks. Improvements in the user interface were recommended to make the tool more intuitive and accessible, with clearer design elements that could facilitate easier navigation and usage. Enhanced error messaging and user feedback, particularly for users with programming experience, were noted as critical for allowing more effective debugging and application refinement.
These findings indicate a significant need for advancements in natural language understanding within Aptly to accommodate a broader array of user instructions and technical demands. The study highlights the essential balance between developing a tool that is accessible for beginners while being sufficiently advanced to support more complex user requirements. Expanding functionalities to include sophisticated web operations and comprehensive multimedia management, along with providing educational examples or templates, could help users better understand and utilize Aptly’s full range of capabilities. This comprehensive feedback not only reflects the current usability challenges faced by Aptly but also provides a strategic roadmap for future enhancements to better serve a diverse user base for tool developers.
IV Related Work
Automatic programming [17, 18], with its rich historical tapestry, has always been about bridging the communication gap between human intent and machine execution. At its core, it embodies the principle of transforming higher-level instructions into forms that a computer can readily understand and act upon. Today’s automatic programming aims to streamline the software development process, enabling developers to emphasize problem definition over intricate implementation details. This evolution promises a future of enhanced collaboration between humans and machines in software creation, making the process more accessible, efficient, and resilient against errors.
Modern automatic programming systems leverage machine learning models to generate code from a higher-level specification or even from natural language descriptions [19]. This has broadened the scope of automatic programming from simple code translation to complex code synthesis, opening up new possibilities for software development [20, 21]. Take, for instance, GitHub Copilot, a coding assistant powered by OpenAI Codex [22]. GitHub Copilot offers contextually relevant code suggestions as you type, effectively automating part of the coding process. Whether you’re learning a new language or building a complex system, Copilot serves as a knowledgeable companion, aiding in writing new code, navigating existing code, and even generating unit tests [23, 24, 25, 26, 27, 28, 29, 30].
Aptly is an attempt to push such boundaries to automatically create block code based on the user’s utterance. Text2App is a similar system based on natural language processing to convert user input into a representation eventually compatible with MIT App Inventor [31]. Such tools are moving toward a direction where anyone can create functional software without any need for traditional coding [32].
V Discussion & Conclusion
Overall, the design and implementation of tools like Aptly should inform how LLMs may apply to other visual programming environments, whether they are block based, flow based, or some other format. Our findings are most applicable to other block based programming environments built on Google Blockly, but the concepts should be translatable to other visual paradigms as well. By designing Aptly to be like Python while maintaining a correspondence to the original visual environment, we are able to leverage the depth of knowledge of LLMs to speed system development. We were also surprised by how well the LLMs were able to transform non-English instructions into working code examples as none of our example prompts were expressed in any language other than English. This suggests that we may see a much lower barriers to entry in the future with regard to computer science education as LLMs gain adoption.
In this work, we described the Aptly language, architecture, and server, and how they form a single system alongside the MIT App Inventor visual programming environment. We demonstrated how large language models are able to synthesize programs in this new language without any training on a new corpus. We further showed how edits to Aptly programs can be translated into corresponding edits in App Inventor projects, enabling Aptly to act as a collaborative programming agent. A small pilot study showed that students were better able to articulate how they would leverage AI systems like Aptly after working with the system, but that there was not a significant shift in their overall interest in learning to program.
The advent of no-code platforms such as Aptly poses both opportunities and challenges for computer science education, which merit careful exploration and discussion. As this trend continues to revolutionize the software industry, it’s becoming increasingly clear that our approach to computer science education may need to evolve to keep pace with these changes [33]. Traditional computer science education has long been centered on manual coding and algorithm design. It emphasizes the need for students to understand the intricacies of programming languages, the efficiency of algorithms, and the design of complex software systems. AI-based code-generation techniques may incentivize students to bypass such productive thinking. Future research on students’ usage of such tools, what they learn while using them [34], and how they feel about using these tools [35] will be active research.
Acknowledgment
We thank the whole MIT App Inventor team who helped throughout the development process, especially our staff and students including Evie Cai, Jacky Chen, Mark Friedman, Maura Kelleher, Jasmine Shone, Joyce Yuan. Thank you also to OpenAI and Drew Houston for their support and resources.
The second listing of Section II-C was generated using Aptly’s generative AI capabilities.